# HeatFormer: A Neural Optimizer for Multiview Human Mesh Recovery

### v1

@@ Abstract
We introduce a novel method for human shape and pose recovery that can fully leverage multiple static views. We target fixed-multiview people monitoring, including elderly care and safety monitoring, in which calibrated cameras can be installed at the corners of a room or an open space but whose configuration may vary depending on the environment. Our key idea is to formulate it as neural optimization. We achieve this with HeatFormer, a neural optimizer that iteratively refines the SMPL parameters given multiview images, which is fundamentally agonistic to the configuration of views. HeatFormer realizes this SMPL parameter estimation as heat map generation and alignment with a novel transformer encoder and decoder. We demonstrate the effectiveness of HeatFormer including its accuracy, robustness to occlusion, and generalizability through an extensive set of experiments. We believe HeatFormer can serve a key role in passive human behavior modeling.
@@ 1 Introduction
Our neural human body shape and pose optimizer, which we refer to as HeatFormer, consists of three key technical contributions. The first is the overall architecture. We build upon the Transformer architecture and derive a neural optimizer that iteratively refines the SMPL parameters given multiview images which is fundamentally agonistic to the number of views and their configurations. The second is the formulation of body shape and pose estimation as heatmap matching. Directly evaluating the error of SMPL parameters in the image space is near impossible as we do not have masks or accurate texture of the person. Keypoint supervision of joints is more realistic, but are hard to estimate with a neural optimizer as they do not offer dense spatial gradients. We turn to the original idea of heatmaps and recast SMPL parameter estimation as heatmap generation and matching. Finally, we derive a novel Transformer encoder that takes in the joint heatmaps and consolidates their spatial coordination into a single heatmap.
@@ 2 Related Work
Table 1: Our method (“Ours”) fully leverages multiview images in the estimation process itself (“Feedback”) to achieve “Occlusion-Robust”ness and “Generalization” to various environments and view configurations. It achieves this by aligning the ”Heatmap” of the view-dependent SMPL estimates with those of the input images. We compare with other multiview methods and demonstrate the effectiveness due to these key strengths.
@@
Our focus is specifically on fully leveraging multiview observations and we leave the use of temporal information for future work. A number of monocular temporal pose estimation methods [21, 8, 47, 22, 50, 48] have also introduced inspiring mechanisms that may apply to instantaneous but multiview HMR. HuMoR [41] integrates optimization in a feed-forward regression framework to ensure foot grounding and temporal consistency of the estimated pose and shape. TCMR [8] extract past and future features to estimate in-between SMPL parameters.
The Transformer architecture has also been leveraged for feed-forward estimation of both non-parametric and parametric models [32, 51, 12, 45, 2, 10, 47, 55, 30, 40]. METRO [32] predicts vertices by relating joints and vertices with a Transformer encoder. PMCE [51] co-learns 3D joints and a SMPL mesh with a Transformer decoder. HMR2.0 [12] extracts deep spatial features with a Transformer encoder and aggregates them in a SMPL query token with a Transformer decoder. Multi-HMR [2] recovers multiple people in a view with a Vision Transformer (ViT) backbone. Our work complements these works with a novel Transformer-based neural optimizer that fully leverages static multiview images. To the best of our knowledge, our method is the first neural optimizer for HMR.
@@ 3 HeatFormer
Figure 2: HeatFormer realizes neural optimization for HMR in which the Transformer encoder-decoder model serves as an unrolled iteration of SMPL fitting to the observed images. It first extracts image features and a heatmap for each view which are aggregated with a novel encoder and input to the decoder. The decoder also takes in heatmaps generated from the current SMPL estimate and through its unrolled inference, iteratively aligns them together.
We achieve occlusion-robust multiview human body shape and pose recovery with a novel transformer which we refer to as HeatFormer. Fig. 2 depicts the overall architecture of HeatFormer. HeatFormer realizes optimization-based HMR in a feed-forward inference in which the Transformer encoder-decoder model serves as an unrolled iteration of SMPL fitting to the observed images. A key idea underlying HeatFormer is to represent and align the body joints as heatmaps. This way, it can be learned robustly end-to-end. HeatFormer forms heatmaps from the current view-dependent SMPL estimates which are iteratively aligned with the input multiview heatmaps through the encoder-decoder inference iteration.
For this, we introduce two key components into the Transformer architecture. One is a heatmap encoder that not only better encodes the 2D joints as visual cues of the body shape and pose in the multiview observations but also realizes effective consolidating of heatmaps across the joints through the self attention. The other is a decoder that cross-attends between the encoder-transformed multiview observations and the current body shape and pose both expressed in heatmaps. The decoder achieves this by taking in the heatmaps formed by the current SMPL estimate as queries. The encoder-decoder inference is iterated multiple times, fully leveraging the images as feedback in the neural optimization to align the SMPL heatmaps without regard to the number of views and their configuration.
@@ 3.1 Preliminaries
We derive a novel Transformer architecture that outputs these SMPL shape and pose parameters for a person captured in multiview images. The input tokens to the encoder correspond to the visual evidence of the human body captured in each view. The encoder learns to consolidate these views which are then cross-attended with the decoder queries. The decoder learns to output a revision to the current SMPL parameter estimate and its queries are computed from the revised SMPL estimate. As a whole, HeatFormer is an unrolled iterative optimization, realized in feed-forward inference, that fits SMPL to the input views.
@@ 3.2 HeatEncoder
The input multiview images are first converted into image features via Vision Transformer (ViT). That is, each view is broken into patches of a regular grid, input to a pre-trained ViT, and the outputs are reassembled into a 2D image feature. We also extract heatmaps for each view. For this, we leverage AdaFuse [54] which is a multiview-aware joint detector. All views are input to AdaFuse which outputs heatmaps for each view, one heatmap for each possible human joint. Now each view is converted into ViT features and multiple ($k$: the number of joints) heatmaps. We next consolidate each set of these heatmaps across different joints for each view.
Fig. 2 shows the overall architecture of HeatEncoder. Given a set of heatmaps, $P\in\mathbb{R}^{k\times H\times W}$, where $H$ and $W$ are the heatmap height and width, respectively, corresponding to one view, HeatEncoder divides the heatmaps into patches $P_{i}\in\mathbb{R}^{k\times h\times w}$, where $i=1,\ldots,\frac{HW}{hw}$ and $(h,w)$ is the patch size. Each patch is flattened and becomes a token $Q\in\mathbb{R}^{k\times\frac{H}{h}\times\frac{W}{w}\times(h\times w)}$, where $k$ is the number of tokens.
@@ 3.3 Decoder As A Neural Optimizer
The outputs of the decoder are reshaped into $N\times C\times(h\times w)$ tensors, where $N$ is the number of views and $C$ is the channel size, and passed through average pooling and an MLP, so that they represent SMPL parameter revisions to the current SMPL estimates, i.e., the difference that should be added to the current estimate. The SMPL parameters are revised and new heatmaps for each view are generated from SMPL instantiated for each view. These heatmaps are used as new queries of the decoder and the revision is computed through cross-attention again. This repeated forward inference realizes neural optimization. In our experiments, we used three iterations, i.e., unrolled forward inferences. As shown with an example in Fig. 3, we observed clear convergence for all cases within these three iterations.
For the initial estimates, we use the SMPL mean parameters for the pose and shape, following previous works such as HMR [20]. The use of mean parameters when computing the global orientation of the body can cause erroneous alignment (e.g., a forced twist in the body). For this, we estimate the global orientation in advance for each view individually with a pre-trained Transformer. The translation can be easily recovered when the cameras are externally calibrated. For uncalibrated cameras, similar to HMR2.0 [12], we use weak-perspective camera parameters regressed from image features for each view. As we cannot leverage multview images for heatmap computation (i.e., AdaFuse [54]) when the cameras are uncalibrated, we directly use the heatmaps recovered from each view independently.
@@ 3.4 Loss Function And Training
Following prior work on HMR, we train our model using losses on the 3D joints, 2D joints, and SMPL parameters, and leverage an adversarial loss. We also add losses on the heatmap and mesh vertices. Our model predicts SMPL parameters $(\theta,\beta)$ and the weak-perspective camera parameters given input multiview images. The 3D joint loss $\mathcal{L}_{\textrm{3D}}$ is computed as the mean-squared error (MSE) of the joints recovered from the estimated SMPL mesh from the ground-truth 3D joints. The 2D joint loss $\mathcal{L}_{\textrm{2D}}$ is computed as the MSE of the view-projected 3D joint estimates from the cropped and normalized ground-truth 2D joints. The SMPL parameter loss $\mathcal{L}_{\textrm{smpl}}$ is computed as the MSE between the SMPL estimates and their annotated (pseudo ground-truth) parameter values in the dataset.
@@
Sometimes the body would try to twist to align the 2D heatmaps. To avoid this, we add vertex losses that measures the MSE of 3D vertices on the SMPL mesh instantiated with the estimated and dataset-annotated SMPL parameters. This loss is computed for each repeated decoder inference. To discourage recovery of unrealistic human pose and shape, we follow HMR [20] and employ an adversarial prior and compute an adversarial loss $\mathcal{L}_{adv}$. The complete loss becomes
@@
where $\mathcal{L}_{\textrm{smpl}}$ is further broken down in to the weighted sum of MSE on the pose $\theta$ and shape $\beta$ parameters with corresponding weights.
@@
Note the HeatFormer is agnostic to the number of views and camera configuration. It can be trained on a number of views different from the number of cameras used at inference time. The decoder iteration can also be varied from 3, but we empirically found the neural optimization to converge in 3 iterations. We experimentally validate all these.
@@ 4 Experiments
We pre-train HeatEncoder on Human3.6M [16] and MPI-INF-3DHP [35] with single-views. We then freeze HeatEncoder and ViT and train the entire HeatFormer on Human3.6M and MPI-INF-3DHP with multiview images. Following previous work [29, 42, 18, 27] on multiview HMR, we train on subjects 1, 5, 6, 7, and 8 and test on subjects 9 and 11 for Human3.6M and train on subjects 1 to 7 and test on subject 8 for MPI-INF-3DHP. We use the metrics MPJPE, PA-MPJPE, and MPVPE to evaluate for Human3.6M dataset. MPJPE/MPVPE is the Mean Per Joint/Vertex Position Error and PA means procrustes aligned. For the MPI-INF-3DHP dataset, we use the metrics MPJPE, PCK, and AUC. Note that we evaluate all metrics after procrustes alignment for MPI-INF-3DHP. PCK (Percentage of Correct Keypoints) represents the percentage of keypoints less than or equal to a predetermined distance from the ground truth. We set the threshold to 150mm following previous works. AUC (Area Under the Curve) represents the area under the PCK curve when changing the PCK threshold in the range of 0-150mm in 5mm increments.
@@
Tab. 2 shows the accuracy of HeatFormer in comparison with state-of-the-art methods for both single-view and multiview methods evaluated on the Human3.6M dataset. The results, of course except ours, are those reported in their respective papers to ensure fairness. All methods are trained on their own combinations of datasets. Our HeatFormer was trained on the training splits of Human3.6M and MPI-INF-3DHP. Ours† uses ground-truth 3D joints to estimate the translation. It is shown as reference to show the effect of translational errors. Note that U-HMR [27] is not published outside of arXiv. Except for this unpublished work, HeatFormer achieves SOTA accuracy and demonstrates the effectiveness of fully leveraging multiview images with a neural optimizer. U-HMR is reported to be trained on Human3.6M and appears to overfit to the data as we show in the following experiments on generalization.
To confirm the effectiveness of neural optimization, we evaluate the inference results after one, two, and three inference steps through the decoder (i.e., iterations) The results shown in Tab. 4 demonstrate that our model quickly converges within three iterations. Fig. 4 shows the visibly comparison results. More qualitative results are shown in the appendix.
@@
U-HMR arXiv [27] 31.0 22.8
HeatFormer (Ours) 30.3 23.2
Reference (Ours†) 28.6 23.1
@@
HeatFormer (Ours) 39.3 99.5 73.1
Reference (Ours†) 36.0 99.5 75.3
@@
iter1 33.9 26.3 41.5 41.8 99.4 71.5
iter2 30.8 23.6 37.3 39.7 99.5 72.8
iter3 30.3 23.2 36.8 39.3 99.5 73.1
@@
To evaluate how well the method generalizes to different scenes and view configurations, we train and test on unseen scenes and also altogether different datasets. We first conduct a cross-dataset evaluation by training on the Human3.6M dataset and testing on the MPI-INF-3DHP dataset. We compare with the multiview method that achieved high-accuracy on Human3.6M, U-HMR [27]. We are not able to conduct this cross-dataset evalutions on other methods as they do not release their code. As shown in Tab. 5, our HeatFormer achieves higher accuracy across all metrics. These results show that our HeatFormer generalizes well to unseen scenes and camera configurations thanks to its neural optimization formulation. On the other hand, the results show that U-HMR is likely overfit to Human3.6M. We also evaluate the accuracy of dataset[16, 35]-trained HeatFormer on the BEHAVE dataset in our results for occlusion evaluation. These results also demonstrate the strong generalizability of our method.
@@
HeatFormer (Ours) 56.0 96.9 62.7
Reference (Ours†) 55.4 96.8 63.1
@@
HeatFormer is robust to occlusions by design as it can fully leverage multiview images through the input tokens, the heatmap consolidation with HeatEncoder, and the heatmap alignment with unrolled decoder inference. We evaluate this robustness to occlusion using the BEHAVE dataset  [3] with object occlusions. In this dataset, we use the 15 body joints except the eye, ear and toe joints of the 25 body joints in openpose [5]. We evaluate the accuracy against the occlusions by objects the person handles and also other body parts with 2 different protocols. One is the evaluation of all joints, the other is the evaluation of joints which have confidence scores higher than $0.3$. These confidence scores are given in the dataset. About 70% of all joints has confidence above $0.3$. We compare the occlusion robustness and also cross-dataset generalization power on the BEHAVE dataset with HMR2.0 [12] as it is a monocular method particularly claimed to be robust to occlusions and also pre-trained on a large set of datasets. We do not compare with U-HMR [27] as we were unable to retrain the method with the available code. Our HeatFormer is trained on the Human3.6M and MPI-INF-3DHP datasets, and HMR2.0 is trained on the Human3.6M, MPI-INF3DHP, COCO [33] and MPII [1] datasets. As the results shown in Tab. 6 clearly show, HeatFormer achieves higher accuracy for both protocols, which demonstrates its robustness to occlusion and also its strength in generalization to different scenes and view configurations.
@@
HeatFormer (Ours) 50.6 31.9 33.8 21.1
Reference (Ours†) 51.8 33.0 34.7 21.7
@@
Tab. 8 shows results on training and testing on different numbers of views. The architecture of HeatFormer allows discrepancy in the number of views used for training and testing and still maintains high accuracy. This flexibility in view configurations, thanks to the Transformer-based neural optimization formulation, is a key strength that would likely significantly adds to the practicality of HeatFormer.

### v3

@@ Abstract
We introduce a novel method for human shape and pose recovery that can fully leverage multiple static views. We target fixed-multiview people monitoring, including elderly care and safety monitoring, in which cameras can be installed at the corners of a room or an open space but whose configuration may vary depending on the environment. Our key idea is to formulate it as neural optimization. We achieve this with HeatFormer, a neural optimizer that iteratively refines the SMPL parameters given multiview images. HeatFormer realizes this SMPL parameter estimation as heatmap generation and alignment with a novel transformer encoder and decoder. Our formulation makes HeatFormer fundamentally agnostic to the number of cameras, their configuration, and calibration. We demonstrate the effectiveness of HeatFormer including its accuracy, robustness to occlusion, and generalizability through an extensive set of experiments. We believe HeatFormer can serve a key role in passive human behavior modeling.
@@ 1 Introduction
Our neural human body shape and pose optimizer, which we refer to as HeatFormer, consists of three key technical contributions. The first is the overall architecture. We build upon the Transformer architecture and derive a neural optimizer that iteratively refines the SMPL parameters given multiview images which is fundamentally agnostic to the number of views and their configurations. The second is the formulation of body shape and pose estimation as heatmap matching. Directly evaluating the error of SMPL parameters in the image space is near impossible as we do not have masks or accurate texture of the person. Keypoint supervision of joints is more realistic, but are hard to estimate with a neural optimizer as they do not offer dense spatial gradients. We turn to the original idea of heatmaps and recast SMPL parameter estimation as heatmap generation and matching. Finally, we derive a novel Transformer encoder that takes in the joint heatmaps and consolidates their spatial coordination into a single heatmap. Our formulation enables application of HeatFormer to uncalibrated cameras. At the same time, when the cameras are calibrated, HeatFormer fully leverages the camera extrinsics.
@@ 2 Related Work
Table 1: Our method (“Ours”) fully leverages multiview images in the estimation process itself (“Feedback”) to achieve “Occlusion-Robust”ness and “Generalization” to various environments and view configurations. It achieves this by aligning the “Heatmap” of the view-dependent SMPL estimates with those of the input images. We compare with other multiview methods and demonstrate the effectiveness due to these key strengths.
@@
Our focus is specifically on fully leveraging multiview observations, and we leave the use of temporal information for future work. A number of monocular temporal pose estimation methods [22, 8, 48, 23, 51, 49] have also introduced inspiring mechanisms that may apply to instantaneous but multiview HMR. HuMoR [42] integrates optimization in a feed-forward regression framework to ensure foot grounding and temporal consistency of the estimated pose and shape. TCMR [8] extract past and future features to estimate in-between SMPL parameters.
The Transformer architecture has also been leveraged for feed-forward estimation of both non-parametric and parametric models [33, 52, 13, 46, 2, 11, 48, 56, 31, 41]. METRO [33] predicts vertices by relating joints and vertices with a Transformer encoder. PMCE [52] co-learns 3D joints and a SMPL mesh with a Transformer decoder. HMR2.0 [13] extracts deep spatial features with a Transformer encoder and aggregates them in a SMPL query token with a Transformer decoder. Multi-HMR [2] recovers multiple people in a view with a Vision Transformer (ViT) [10] backbone. Our work complements these works with a novel Transformer-based neural optimizer that fully leverages static multiview images. To the best of our knowledge, our method is the first neural optimizer for HMR.
@@ 3 HeatFormer
Figure 2: HeatFormer realizes neural optimization for HMR in which the Transformer encoder-decoder model serves as an unrolled iteration of SMPL fitting to the observed images. It first extracts image features and a heatmap for each view which are aggregated with a novel encoder and input to the decoder. The decoder also takes in heatmaps generated from the current SMPL estimate and, through its unrolled inference, iteratively aligns them together.
We achieve occlusion-robust multiview human body shape and pose recovery with a novel transformer which we refer to as HeatFormer. Fig. 2 depicts the overall architecture of HeatFormer. HeatFormer realizes optimization-based HMR in a feed-forward inference in which the Transformer encoder-decoder model serves as an unrolled iteration of SMPL fitting to the observed images. A key idea underlying HeatFormer is to represent and align the body joints as heatmaps. HeatFormer forms heatmaps from the current view-dependent SMPL estimates which are iteratively aligned with the input multiview heatmaps through the encoder-decoder inference iteration.
For this, we introduce two key components into the Transformer architecture. One is a heatmap encoder that not only better encodes the 2D joints as visual cues of the body shape and pose in the multiview observations but also realizes effective consolidation of heatmaps across the joints through self-attention. The other is a decoder that cross-attends between the encoder-transformed multiview observations and the current body shape and pose both expressed in heatmaps. The decoder achieves this by taking in the heatmaps formed by the current SMPL estimate as queries. The encoder-decoder inference is iterated multiple times, fully leveraging the images as feedback in the neural optimization to align the SMPL heatmaps without regard to the number of views and their configuration.
@@ 3.1 Preliminaries
We derive a novel Transformer architecture that outputs these SMPL shape and pose parameters for a person captured in multiview images. The input tokens to the encoder correspond to heatmaps of the human body captured in each view. The encoder learns to consolidate heatmaps in each view which are then cross-attended with the decoder queries. The decoder learns to output a revision to the current SMPL parameter estimate and its queries are computed from the revised SMPL estimate. As a whole, HeatFormer is an unrolled iterative optimization, realized in feed-forward inference, that fits SMPL to the input views.
@@ 3.2 HeatEncoder
The input multiview images are first converted into image features via Vision Transformer (ViT) [10]. That is, each view is broken into patches of a regular grid, input to a pre-trained ViT, and the outputs are reassembled into a 2D image feature. We also extract heatmaps for each view. For this, we leverage AdaFuse [55] which is a multiview-aware joint detector. All views are input to AdaFuse which outputs heatmaps for each view, one heatmap for each possible human joint. Now each view is converted into ViT features and multiple ($k$: the number of joints) heatmaps. We next consolidate each set of these heatmaps across different joints for each view.
Fig. 2 shows the overall architecture of HeatEncoder. Given a set of heatmaps, $P\in\mathbb{R}^{k\times H\times W}$, where $H$ and $W$ are the heatmap height and width, respectively, corresponding to one view, HeatEncoder divides the heatmaps into patches $P_{i}\in\mathbb{R}^{k\times h\times w}$, where $i=1,\ldots,\frac{HW}{hw}$ and $(h,w)$ are the patch size. Each patch is flattened and becomes a token $Q\in\mathbb{R}^{k\times\frac{H}{h}\times\frac{W}{w}\times(h\times w)}$, where $k$ is the number of tokens.
@@ 3.3 Decoder As A Neural Optimizer
The outputs of the decoder are reshaped into $N\times C\times(h\times w)$ tensors, where $N$ is the number of views and $C$ is the channel size, and passed through average pooling and an MLP, so that they represent SMPL parameter revisions to the current SMPL estimates, i.e., the difference that should be added to the current estimate. The SMPL parameters are revised and new heatmaps for each view are generated from SMPL instantiated for each view. These heatmaps are used as new queries of the decoder and the revision is computed through cross-attention again. This repeated forward inference realizes neural optimization. In our experiments, we used three iterations, i.e., unrolled forward inferences. As shown with an example in Fig. 3, we observed clear convergence for all cases within these three or four iterations.
For the initial estimates, we use the SMPL mean parameters for the pose and shape, following previous works such as HMR [21]. The use of mean parameters when computing the global orientation of the body can cause erroneous alignment (e.g., a forced twist in the body). For this, we estimate the global orientation in advance for each view individually with a pre-trained Transformer. The translation can be easily recovered when the cameras are externally calibrated. HeatFormer can be also applied without camera calibration. The camera extrinsics are only used for opportunistic sanitization of heatmaps with AdaFuse and query heatmap projection. These steps can be either skipped or instead estimated, respectively, when the camera extrinsics are unknown. For the latter, when using uncalibrated cameras, we can use the weak-perspective camera parameters regressed from image features for each view similar to HMR2.0 [13]. Please see the appendix for further details.
@@ 3.4 Loss Function And Training
Following prior work on HMR, we train our model using losses on the 3D joints, 2D joints, and SMPL parameters, and leverage an adversarial loss. We also add losses on the heatmap and mesh vertices. Our model predicts SMPL parameters $(\theta,\beta)$ and the weak-perspective camera parameters given input multiview images. The 3D joint loss $\mathcal{L}_{\textrm{3D}}$ is computed as the mean-squared error (MSE) of the joints recovered from the estimated SMPL mesh from the ground-truth 3D joints. The 2D joint loss $\mathcal{L}_{\textrm{2D}}$ is computed as the MSE of the view-projected 3D joint estimates from the cropped and normalized ground-truth 2D joints.
The SMPL parameter loss $\mathcal{L}_{\textrm{smpl}}$ is computed as the MSE between the SMPL estimates and their annotated (pseudo ground-truth) parameter values in the dataset.
@@
Sometimes the body would try to twist to align the 2D heatmaps. We avoid this with a vertex loss that measures the MSE of 3D vertices on the SMPL mesh instantiated with the estimated and dataset-annotated SMPL parameters. This loss is computed for each repeated decoder inference.
We follow HMR [21] and employ an adversarial prior and compute an adversarial loss $\mathcal{L}_{adv}$ to discourage unrealistic human poses. The complete loss becomes
@@
where $\mathcal{L}_{\textrm{smpl}}$ is a weighted sum of MSE on the pose $\theta$ and shape $\beta$ parameters with corresponding weights.
@@
Note that HeatFormer is agnostic to the number of views and camera configuration. It can be trained on a number of views different from the number of cameras used at inference time. The decoder iteration can also be varied, but we empirically found that the neural optimization converges in 3 or 4 iterations. We experimentally validate all these.
@@ 4 Experiments
We pre-train HeatEncoder on Human3.6M [17] and MPI-INF-3DHP [36] with single-view. We then freeze HeatEncoder and ViT and train the entire HeatFormer on Human3.6M and MPI-INF-3DHP with multiview images. Following previous work [30, 43, 19, 28] on multiview HMR, we train on subjects 1, 5, 6, 7, and 8 and test on subjects 9 and 11 for Human3.6M and train on subjects 1 to 7 and test on subject 8 for MPI-INF-3DHP. We use the metrics MPJPE, PA-MPJPE, and MPVPE to evaluate for Human3.6M dataset. MPJPE/MPVPE is the Mean Per Joint/Vertex Position Error and PA means procrustes aligned. For the MPI-INF-3DHP dataset, we use the metrics MPJPE, PCK, and AUC. Note that we evaluate all metrics after procrustes alignment for MPI-INF-3DHP. PCK (Percentage of Correct Keypoints) represents the percentage of keypoints less than or equal to a predetermined distance from the ground truth. We set the threshold to 150mm following previous works. AUC (Area Under the Curve) represents the area under the PCK curve when changing the PCK threshold in the range of 0-150mm in 5mm increments.
@@
Tab. 2 shows the accuracy of HeatFormer in comparison with state-of-the-art methods for both single-view and multiview methods evaluated on the Human3.6M dataset. The results, of course except ours, are those reported in their respective papers to ensure fairness. All methods are trained on their own combinations of datasets. Our HeatFormer was trained on the training splits of Human3.6M and MPI-INF-3DHP. Ours† uses ground-truth 3D joints to estimate the translation. It is shown as reference to show the effect of translational errors. HeatFormer achieves SOTA accuracy and demonstrates the effectiveness of fully leveraging multiview images with a neural optimizer.
To confirm the effectiveness of neural optimization, we evaluate the inference results after one, two, and three inference steps through the decoder of a 3-iter model. The results shown in Tab. 4 show that our model quickly converges. Fig. 4 shows visual comparisons. Please see the appendix for more qualitative results.
@@
U-HMR [28] 31.0 22.8
HeatFormer (Ours iter3) 30.7 23.3
HeatFormer (Ours iter4) 29.5 22.4
Reference (Ours† iter3) 28.6 23.1
@@
HeatFormer (Ours iter3) 39.8 99.5 72.8
HeatFormer (Ours iter4) 40.6 99.5 72.3
Reference (Ours† iter3) 36.0 99.5 75.3
@@
iter1 34.9 26.2 41.9 44.6 99.3 70.0
iter2 31.2 23.6 37.5 40.8 99.5 72.1
iter3 30.7 23.3 37.0 39.8 99.5 72.8
@@
To evaluate how well the method generalizes to different scenes and view configurations, we train and test on unseen scenes and also altogether different datasets. We first conduct a cross-dataset evaluation by training on the Human3.6M dataset and testing on the MPI-INF-3DHP dataset. We compare with the multiview method that achieved high-accuracy on Human3.6M, U-HMR [28]. As shown in Tab. 5, our HeatFormer achieves higher accuracy across all metrics. These results show that our HeatFormer generalizes well to unseen scenes and camera configurations thanks to its neural optimization formulation. On the other hand, the results show that U-HMR is likely overfit to Human3.6M. We also evaluate the accuracy of HeatFormer trained on Human3.6M and MPI-INF-3DHP [17, 36] when applied to the BEHAVE dataset in our results for occlusion evaluation. These results also demonstrate the strong generalizability of our method.
@@
HeatFormer (Ours iter3) 56.0 96.9 62.7
Reference (Ours† iter3) 55.4 96.8 63.1
@@
HeatFormer is robust to occlusions by design as it can fully leverage multiview images through the input tokens, the heatmap consolidation with HeatEncoder, and the heatmap alignment with unrolled decoder inference. We evaluate this robustness to occlusion using the BEHAVE dataset [3] with object occlusions. In this dataset, we use the 15 body joints except the eye, ear and toe joints of the 25 body joints in openpose [5]. We evaluate the accuracy against the occlusions by objects the person handles and also other body parts with 2 different protocols. One is the evaluation of all joints, the other is the evaluation of joints which have confidence scores higher than $0.3$. These confidence scores are given in the dataset. About 70% of all joints has confidence above $0.3$. We compare the occlusion robustness and also cross-dataset generalization power on the BEHAVE dataset with HMR2.0 [13] as it is a monocular method particularly claimed to be robust to occlusions and also pre-trained on a large set of datasets and multiview refinement methods [26, 45]. Our HeatFormer is trained on the Human3.6M and MPI-INF-3DHP datasets, and HMR2.0 is trained on the Human3.6M, MPI-INF3DHP, COCO [34] and MPII [1] datasets. As the results in Tab. 6 clearly show, HeatFormer achieves high accuracy for both protocols, which demonstrates its robustness to occlusion and also its strength in generalization to different scenes and view configurations.
@@
ProHMR [26] 111.2 55.7 74.7 37.1
HMR2.0a+scoreHMR [45] 72.9 30.9 49.0 20.1
HMR2.0b+ScoreHMR [45] 86.0 35.3 58.0 22.8
HeatFormer (Ours iter3) 51.1 32.2 34.2 21.2
HeatFormer (Ours iter4) 48.9 31.8 32.6 20.9
Reference (Ours† iter3) 51.8 33.0 34.7 21.7
@@
Tab. 8 shows results on training and testing on different numbers of views. The architecture of HeatFormer allows discrepancy in the number of views used for training and testing and still maintains high accuracy. This flexibility in view configurations, thanks to the Transformer-based neural optimization formulation, is a key strength that likely significantly adds to the practicality of HeatFormer.